Find a linear equation whose graph is the straight line with the given properties.
Through (5,45) with slope 9
y(x)=

Answer by rothauserc(4718)

Use the point slope form of a line
:
y = 9x +b, where b is the y-axis intercept
:
Now use the point (5, 45)
:
45 = 9*5 +b
:
b = 45 - 45 = 0
:
*******
y = 9x
*******
: